A survey of 1000 students at a large university asked students about their view on capital punishment. Half of the students surveyed were men and half were women. Of the men surveyed, 280 were opposed to capital punishment. Of the women surveyed, 320 were opposed.
(a) How many students in the sample were in favor of capital punishment? 400
(b) How many of those in favor of capital punishment were women?

Answer by ikleyn(52781)

.

1000 students = 500 men + 500 women.


(a)  How many students in the sample were in favor of capital punishment = (500-280) + (500-320) = 220 + 180 = 400.    ANSWER


(b)  How many of those in favor of capital punishment were women = (500-320) = 180.    ANSWER

Solved.

As simple as 2 + 3 = 5.

This task is just a test for the presence of common sense in the head of the testee.